jagr2713 (jagr2713):

Let f(x) = x2 - 81. Find f-1(x). ±square root of the quantity x plus 81 9square root of x 1 over quantity x squared minus 81 x squared over 81

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

jagr2713 (jagr2713):

yea i did that but i got confused

OpenStudy (anonymous):

I've seen you attempt harder problems, show me where you went wrong :P

jagr2713 (jagr2713):

this is where i stop at y=5x+12 x=5y+12 x-12/5=5y/5 x-12/5=fx

OpenStudy (anonymous):

Wrong problem? y =5x+12?

jagr2713 (jagr2713):

opps wrong one i got the answer to that one i mean i really didn't start the question i am asking you

OpenStudy (anonymous):

Lol well, follow the steps I gave you, and we can see how you do :)

OpenStudy (anonymous):

\[f(x) = x^-81 \implies y = x^2-81\]

OpenStudy (anonymous):

Do the next step :D

jagr2713 (jagr2713):

is it C

OpenStudy (anonymous):

Jagr I don't understand, it's just simple algebra, I've seen you answer much tougher questions..mhm, and no actually do the problem instead of guessing. Just follow the steps I gave you, I'll help you.

OpenStudy (anonymous):

So we have y=x^2-81, next step?

jagr2713 (jagr2713):

x-81=y^2

OpenStudy (anonymous):

No, that's not right, you just replaced the square from x to y, not the variable itself.

jagr2713 (jagr2713):

THIS is CONFUSING. ok so was i even close

OpenStudy (anonymous):

\[y=x^2-81, ~~~ x=y^2-81\] that's all you had to do, don't over complicate it.

jagr2713 (jagr2713):

oh ok

OpenStudy (anonymous):

Now solve for y

jagr2713 (jagr2713):

x squared over 81

jagr2713 (jagr2713):

do you square root 81 to 9

OpenStudy (anonymous):

\[y = \pm \sqrt{x+81}\]

OpenStudy (anonymous):

Now just change the y back to f^(-1)

jagr2713 (jagr2713):

so its A

jagr2713 (jagr2713):

@iambatman is it A

OpenStudy (anonymous):

Yes it is A.